OpenStudy (anonymous):

It takes a pump 3 hours to fill a cylindrical barrel. How long would it take the pump to fill a barrel with the same height but twice the radius? Please show the work.

OpenStudy (anonymous):

bet it would be 4 times as long. lets check

OpenStudy (anonymous):

the answer is we hours so you are right i jsut want to see the work

OpenStudy (anonymous):

since you know neither height nor radius it doesn't matter so we can make up numbers. suppose the radius of the first barrel is 1 and the height is h. then the volume is \[\pi r^2h = \pi h\] now if the radius doubles to be 2, then the volume is \[ \pi \times 2^2 \times h = 4\pi h\] so the volume is increased by a factor of 4

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

therefore it will take 4 times as long

OpenStudy (anonymous):

yw